# Yantao Linormal-$\diamondsuit$   Jianbing Zhangnormal-$\diamondsuit$   Zhiyong Wunormal-$\heartsuit$

### 17 Jan 2024 v1

@@ B.1 MiniWob
Figure 6(a) illustrates the difference between static and dynamic layout tasks. Static layout tasks have fixed element positions during training and testing, while dynamic layout tasks display varying interfaces and element positions. Figure 6(b) provides examples of SeeClick’s interaction with MiniWob. SeeClick relies solely on the interface screenshot for arithmetic, reasoning, etc.
B.2 AITW
@@
Figure 7 provides an example of SeeClick when completing an AITW task. Predictions marked in red below indicate that they were computed as incorrect in AITW. However, some errors occur because the current step’s answer is not unique. For example in step 5, the model’s predicted input "DuckDuckGo Privacy Browser" is also a potentially correct action.
B.3 Mind2Web
@@
Figure 8 shows several examples of SeeClick on the real-world website benchmark Mind2Web. SeeClick can comprehend instructions and click on the correct elements within complex interfaces.
@@
Figure 6: Example episodes of SeeClick on MiniWob. The model’s prediction output is below the screenshot, with action_type 4 indicating clicking and action_type 3 indicating typing.
Figure 7: Example episodes of SeeClick on AITW. The model’s prediction output is below the screenshot, with action_type 4 indicating clicking, action_type 3 indicating typing and action_type 6 indicating PRESS HOME. For failed steps, we mark the predictions in red and provide the groundtruth action in green.
Figure 8: Example episodes of SeeClick on Mind2Web. The model’s prediction output is below the screenshot, with action_type 4 indicating clicking and action_type 3 indicating typing. For failed steps, we mark the predictions in red and provide the groundtruth elements with the green bounding box.

### 23 Feb 2024 v2

@@ C.5 Case Study
MiniWob Figure 11(a) illustrates the difference between static and dynamic layout tasks. Static layout tasks have fixed element positions during training and testing, while dynamic layout tasks display varying interfaces and element positions with instructions, further challenging the agent’s ability to accurately locate the target. Figure 11(b) provides examples of SeeClick’s interaction with MiniWob. SeeClick relies solely on the interface screenshot for arithmetic, reasoning, etc.
@@
AITW Figure 12 provides SeeClick’s operations on AITW. Predictions marked in red below indicate that they were computed as incorrect in AITW. Some errors occur because the current step’s answer is not unique. For example in step 5, the model’s predicted input "DuckDuckGo Privacy Browser" is also a potentially correct action.
@@
Mind2Web Figure 13 shows several examples of SeeClick on the real-world website benchmark Mind2Web. SeeClick can comprehend instructions and click on the correct elements within complex interfaces.
@@
Figure 11: Example episodes of SeeClick on MiniWob. The model’s prediction output is below the screenshot, with action_type 4 indicating a click and action_type 3 indicating typing.
Figure 12: Example episodes of SeeClick on AITW. The model’s prediction output is below the screenshot, with action_type 4 indicating a click, action_type 3 indicating typing and action_type 6 indicating PRESS HOME. Steps with the red prediction and green reference indicate a failed step.
Figure 13: Example episodes of SeeClick on Mind2Web. The model’s prediction output is below the screenshot, with action_type 4 indicating a click and action_type 3 indicating typing. Steps with the red prediction and green reference bounding box indicate a failed step.